The work done by the electric motor during 30 revolutions of its rotor It is 1266 J. Find the torque developed by the motor.


1
Expert's answer
2021-04-13T06:38:36-0400

The work "W" relates to torque "\tau" as follows (see https://brilliant.org/wiki/rotational-kinetic-energy-work-kinetic-theorem/):


"W = \tau \theta"

where "\theta" is the angle in radians by which the motor revolved. During the 1 revolution the motor covers "2\pi" radians. Thus, during 30 revolutions it covers:


"\theta = 30\cdot 2\pi\space = 60\pi \space"

Thus, the troque is:


"\tau = \dfrac{W}{\theta}\\
\tau = \dfrac{1266J}{60\pi \space } \approx 6.7N\cdot m"

Answer. 6.7 Nm.